OpenStudy (anonymous):

A quadrilateral has angles that measure 44°, 89°, and 127°. What is the measure of the fourth angle?

OpenStudy (anonymous):

all angles in a quadrilateral add up to 360°

OpenStudy (anonymous):

x + 44 + 89 + 127 = 360

OpenStudy (anonymous):

how do i ged the aswer?

OpenStudy (defiance):

44 + 89 + 129=262 They total 360 360 - 262 = 98

OpenStudy (defiance):

You add them, then you subtract it by 360

OpenStudy (defiance):

You have 360 degrees in total, to 360 - <The total of other angles> makes the answer of the last

OpenStudy (anonymous):

i have a nother quastion you will ged to many medalls